t-u=1
t+v=2
v-u=7
Trying to help my daughter solve and am completely lost. I appreciate your help !

Answer by stanbon(75887)

t-u=1
t+v=2
v-u=7
----------
Express u and v in terms of "t".
u = t-1
v = 2-t
----
Substitute into the 3rd equation so you can solve for "t":
(2-t) - (t-1) = 7
2-t-t+1 = 7
-2t + 3 = 7
-2t = 4
t = -2
--------
Since u = t-1, u = -2-1 = -3
Since v = 2-t, v = 2--2 = 2+2 = 4
---
Final answer:
t = -2 ; u = -3 ; v = 4
============================
Check the answers in the original equations:
t-u=1::::-2--3 = 1 ; -2+3 = 1 ; 1=1
t+v=2::::-2+4 = 2 ; 2=2
v-u=7:::: 4--3 = 7 ; 4+3 = 7 ; 7=7
====================================
Cheers,
Stan H.